grace told her mum that the sum of there age is half a century and the difference between them is 2 dozen. is grace a teenager?
i tryed to work out the answer by an equation but im not sure if its right

Answer by Edwin McCravy(20060)

Grace is x
Mum is 50-x

Mum's age minus Grace's age = 24

 (50-x) - x = 24
 50 - x - x = 24
    50 - 2x = 24
        -2x = -26
          x = 13

Yes, she's just barely a teenager.

Edwin